2.  Show that

  (a) var(X – Y) = var(X) + var(Y) – 2cov(X,Y)                                                                                  (b) cov(X,aY + b) = acov(X,Y) if a and are constants.         


1
Expert's answer
2021-07-16T04:02:50-0400

(a)

"Var(X-Y)=E[(X-Y)^2]-(E[X-Y])^2"

"=E[X^2-2XY+Y^2]-(\mu_X-\mu_Y)^2"

"=E[X^2]-2E[XY]+E[Y^2]-\mu_X^2+2\mu_X\mu_Y-\mu_Y^2"


"=(E[X^2]-\mu_X^2)+(E[Y^2]-\mu_Y^2)-2(E[XY]-\mu_X\mu_Y)"

"=Var(X)+Var(Y)-2Cov(X, Y)"

(b)


"Cov(X, aY+b)=E((X-E(X))(aY+b-E(aY+b)))"

"=E((X-E(X))(aY+b-aE(Y)-b))"

"=E((X-E(X))(aY-aE(Y)))"

"=aE((X-E(X))(Y-E(Y)))"

"=aCov(X, Y),"

"a" and "b" are constants.